A car drives around a flat 119 m radius circular track at 20 m/s. Assume that this speed is the maximum speed that the car can have without out "skidding out" of the curve.The car moves into the next curve. The radius of this new curve is twice as great as the previous. Assume the coefficient of static friction has not changed. Calculate the car's maximum speed in this curve.
m/s
Answer #1

max speed a car can have without skidding

v^2 = u r g

here square of velocity is directly proportional to radius

so

(v2/v1) ^2 = (R2 /r1)

(v2 / 20) ^2 = 119* 2 / 119

v2 = 28.284 m/s

======

Comment in case any doubt, will reply for sure Goodluck